OpenStudy (lisa123):

Factor the algebraic expression below in terms of a single trigonometric function.

OpenStudy (lisa123):

cos x - sin ^2x - 1

OpenStudy (lisa123):

Im a bit confused on how solve this problem. Please help!

OpenStudy (anonymous):

replace \(\sin^2(x)\) by \(1-\cos^2(x)\) don't forget the parentheses

Directrix (directrix):

cos x - sin ^2x - 1 = cos x - (1 - cos² x) - 1 = What next? @lisa123

OpenStudy (lisa123):

(cosx)( 1+cosx)-2 ????? @Directrix

OpenStudy (lisa123):

No sorry I meant cos x + cos^2x -2

Directrix (directrix):

cos x - (1 - cos² x) - 1 = cos x - 1 + cos² x - 1 = ? note the use of the distributive property to simplify - (1 - cos² x) @lisa123

OpenStudy (lisa123):

-2 + cos^2x+ cosx

Directrix (directrix):

Question: do you see how to factor y ² + y - 2 ? Because I think this is the same type factoring with y as cox x.

Directrix (directrix):

So, factor this: y ² + y - 2

OpenStudy (lisa123):

y(y+1) -2???

Directrix (directrix):

No. y ² + y - 2 = (y + 2 ) * (y - 1)

Directrix (directrix):

Now, factor this the same way: cos^2x+ cos x - 2

Directrix (directrix):

( cos x + 2 ) * (cos x - 1) @lisa123

OpenStudy (lisa123):

Ok I see thank you this helped a lot!

Directrix (directrix):

You are welcome.